Five yrs ago,a mother was twice as old as his son.in 6yrs time,d sum of their ages will b 82.find their present ages.
Answer by psbhowmick(878)

Let the current age of son be x years and of mother be y years.

5 years ago:
y-5=2(x-5)
y=2x-5
2x-y=5

6 years down:
(x+6)+(y+6)=82
x+y=70

x=25
Y=45